Find the slope of the tangent line to the ellipse x^2/9+y^2/4=1 at the point (x,y)
slope =?
Are there any points where the slope is not defined? (Enter them as comma-separated ordered-pairs, e.g., (1,3), (-2,5). Enter none if there are no such points.)
slope is undefined at =?

Answer by Alan3354(69443)

Find the slope of the tangent line to the ellipse x^2/9+y^2/4=1 at the point (x,y)
-------
x^2/9+y^2/4=1
4x^2 + 9y^2 = 36
8xdx + 18ydy = 0
4xdx = -9ydy
dy/dx = -4x/9y
======================
Are there any points where the slope is not defined? (Enter them as comma-separated ordered-pairs, e.g., (1,3), (-2,5). Enter none if there are no such points.)
Undefined where the DEN is zero.
--> y = 0
x^2/9 = 1
(-3,0) and (3,0)